OpenStudy (anonymous):

−2(x + 3) = −2(x + 1) − 4

OpenStudy (anonymous):

you can't solve this

OpenStudy (anonymous):

the answer is 0

OpenStudy (anonymous):

just FOIL it, and you will see

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

-2x-6=-2x-2-4

OpenStudy (anonymous):

you are welcome!

OpenStudy (anonymous):

3(x + 1) = −2(x − 1) + 6 can you help me with this math problem please

OpenStudy (anonymous):

3x+3=-2x+2+6

OpenStudy (anonymous):

can you do this now?

OpenStudy (anonymous):

x's in one side and numbers in the other side

OpenStudy (anonymous):

its still sort of confusing